Presentation is loading. Please wait.

Presentation is loading. Please wait.

SPS LS1 Planning meeting (N018) 31/07/13

Similar presentations


Presentation on theme: "SPS LS1 Planning meeting (N018) 31/07/13"— Presentation transcript:

1 SPS LS1 Planning meeting (N018) 31/07/13
David Mcfarlane

2 Topics Lifts and Monte-charges. Removal of equipment in LSS1+
Unclaimed cables in BA1 ECX5 Works around the CCC (Surface) Shielding Wall between TT40 and TJ8 Planning. False floors open in the BAs E.C.Rs. Access rights and safety equipment Important Information. A.O.B

3 Lift Maintenance July/August

4 Removal of equipment in LSS1+
The work is proceeding well All 3 dumps have been removed and the dose rate was a lot lower than foreseen. There are status update meetings every Monday morning detailing the works planned for the coming week.

5 LSS1+ planning for week 30 29 July – 02 August
Monday morning Remove TBSJB11995 Monday afternoon Remove remaining BLM  Tuesday – Wednesday Disconnect vacuum up to 11870 Prepare QDA and QFA Thursday Remove the (not so radioactive) QFA11810 Remove QDA11910 MDVA11904 Start labelling cable positions

6 Unclaimed cables in BA1 Je suis en train de traiter chacun d’entre eux, mais après avoir fait un petit point sur la situation, il s’avère que certains des câbles n’ont été checkés par aucun utilisateur. Sur les 1112 câbles concernés par la campagne, j’ai reçu 830 validations. Il me reste donc 230 câbles « orphelins ». Pour certains, je n’ai pas eu de retour de la part des utilisateurs, merci de me le faire suivre urgemment. Pour d’autres, ils devaient faire des recherches plus approfondies afin de s’assurer de la décision à prendre, merci de me tenir informé si les choses ont évolué. Je vous demanderais s’il vous plait, de regarder à nouveau la liste en attachement, et me faire savoir si certains des câbles seraient passés à travers « les mailles du filet ». Dans le cas ou personne serai en mesure de se prononcer sur les câbles restant, je demanderai à la coordination et responsables du complexe SPS, de me donner leur aval pour la suppression de ces derniers. En vous remerciant par avance de votre coopération, Salutations Guillaume Gros

7 ECX5

8 Works around the CCC Installation of new Transformers
Creation of the new cable trench from the CCC to BA3 Works inside Building 874 (this includes works on the outside of the building. Access required around CCC for HVAC replacement works Installation of PIPES trench between CCC and 774 Installation of CABLE trench between CCC and 774 Installation of WASTE WATER trench between CCC and 774 (near Jura) Waste water pump and pipe work (Seleve side of CCC) Changing of SPS ring-main valves Fibre Optics Works (Jura side of CCC)

9 Planning for the next few months

10

11 Shielding Wall between TT40 and TJ8
Dear all, We are re-measuring all the links between the different accelerators and we got trouble with one shielding wall between TT40 and TJ8. In order to pass the horizontal measurements from the SPS through TT40 and TI8 towards the LHC we would need to pass the shielding wall in TT40. This link has not been done since the installation in 2002 and it would be a unique opportunity to do this and to modify the shielding to allow an easier opening in the future. I have met Sylvain this afternoon to have a look on site. In a first approach I`m asking for the partial opening of the shielding to allow the line of sight to pass. I will make the EDH request right now. Does anybody, in particular RP has any objections for the opening? In a second approach one needs to see if the shielding wall can be closed in an way that will allow an easier opening in the future…. Please find enclosed a description of the needed modification. Best regards, Patrick

12

13 SPS LS1 schedule Beam on EPC Commissioning tests
Beam off : 16/02/13 Beam on: 13/10/14 Open Access : 27/03/13 Close Access : 27/06/14 Beam on Cool down/RP survey and Magnet tests Cool down NO ACCESS Fibres 18 KV Kicker Survey Civil Engineering in TT10 RF upgrade works Survey RF upgrade works EPC works Cables Survey Survey Kicker Cables Survey EPC works Survey Survey Maintenance General Maintenance General Magnets Magnets Magnets Magnets Kicker Magnets Magnets Cables Magnets 18 KV Magnets General maintenance General maintenance Crab Cavities General Maintenance Cooling tower 863 General Maintenance Survey Vacuum Ring-main Ring-main Ring-main Vacuum Ring-main Septa works Ring-main Vacuum Ring-main Fibres Septa works Ring-main Prep work in LSS1 18 KV General Maintenance Maintenance General Vacuum Vacuum RF Upgrade works Cables Survey Horizontal Vertical smoothing Cool down Limited ACCESS Fibres TT20 pipework 18 KV CV maintenance CV maintenance UA works Civil Engineering Irradiated Cables 18 KV Fibres CV maintenance Fibres Cables Cables Cables 18 KV Cables EL Studies Civil Engineering TT10 Consolidation Cables replacement Septa Civil Engineering Civil Engineering CV studies RF Upgrade works Crab Cavities Civil Engineering Civil Engineering RF Upgrade works Re-install equipment removed for cable works RAL Magnets CV works Fibres Cables Magnets Magnets Magnets Magnets Magnets Magnets Magnets EPC Commissioning tests Vacuum Cold check out & Start-up DSO tests/HV tests Protons to the North Area Protons to LHC Ions to NA61 CV Controls upgrade Re-route Cabling

14 False floors lifted in Surface BAs
Below is a table showing the dates that EL will be pulling cables at each point of the SPS. During these periods large sections of the false floor in these surface buildings will be lifted. Any works in the BA’s that involve lifting of the false floors MUST be reported to the Coordination team BAs Cabling pulling start Cable pulling end BA1 (TS1- works) BA1 (Irradiated cable works) BA1 (Fibre Optics) 29/04/2013 08/10/2013 08/04/2013 02/08/2013 14/02/2014 24/04/2013 BA2 06/01/2014 11/03/2014 BA3 BA3 (Fibre Optics) 27/03/2013 22/04/2014 31/05/2013 06/05/2014 BA4 27/05/2013 12/07/2013 BA5 BA5 (Fibre Optics) 30/09/2013 24/05/3013 11/10/2013 BA6 BA6 (Fibre Optics) 13/5/2013 18/10/2013 24/05/2013 BA7 16/10/2013 07/11/2013 If you require access to racks or other equipment in the surface buildings during these periods you must check with planning first.

15 Engineering Change Requests
Templates and Guidelines for Engineering Change Requests: EDMS: SPS ECR template (in DFS): https://dfsweb.web.cern.ch/dfsweb/Services/DFS/DFSBrowser.aspx/Applications/CERN/Workgroup%20Templates/SPS/ECR-Template-SPS.dotx

16 Current ECR’s for LS1 https://edms.cern.ch/nav/P:CERN :V0/P:CERN :V0/TAB3 Removal of AEWB Pick-up from position to 31502 (EDMS No ) RELEASED Removal of a BPWA Pick-up from position to 31101 (EDMS No ) RELEASED Installation of a new Wideband Kickers system in position 32173 (EDMS No ) RELEASED Installation of a new HBW Damper Pick-up in place of BPWA & AEWB 31932 (EDMS No ) RELEASED Replacement of 4 old BPA and 2 BPD by 6 BPCE in LSS1 & LSS2 (EDMS No ) RELEASED Reinstallation of pump and gauges (Pirani + Penning) in (vacuum sector 150) (EDMS No ) RELEASED Modification LSS1 sectorization (new VVSF + displacement VVSB 11772) (EDMS No v.2) RELEASED Installation of 3 BCT in LSS5 for North Area protection during Ions operations in SPS. (EDMS No ) UNDER IMPLEMENTATION

17 Current ECR’s for LS1 https://edms.cern.ch/nav/P:CERN :V0/P:CERN :V0/TAB3 Transfer of the Fast Beam Current Transformers (BCTFR.31451) for the SPS circulating beam from LSS3 to LSS5. (EDMS No ) RELEASED SPS Beam Position Monitoring Prototyping Setup in LSS4. (EDMS No ) RELEASED Installation de 4 moniteurs de position (2 BPH, 2 BPV), dédiés au système Damper. (EDMS No ) RELEASED Installation d'un nouveau moniteur de position de type BPCR en LSS2 au (EDMS No ) RELEASED Displacement of Door PPG2145 separating the LHC and the SPS in the TI2 tunnel. (EDMS No ) RELEASED Interlocking of Injection Tunnel upstream areas with LHC powering. (EDMS No ) RELEASED Move of BPH to BPH (EDMS No ) RELEASED Optical Fibre Racks. (EDMS No ) RELEASED

18 Current ECR’s for LS1 https://edms.cern.ch/nav/P:CERN :V0/P:CERN :V0/TAB3 Removal of old fixed TMP vacuum groups in (EDMS No ) RELEASED BYBPM racks under the SPS dipole magnets as part of the Upgrade of Beam Position Monitors electronics (MOPOS). (EDMS No ) UNDER IMPLEMENTATION Installation de Coffrets d'Alimentation en LSS1 et TT10. (EDMS No ) RELEASED Installation d'un rack de jonction RX11695=TS1 au BA1, afin de limiter les longueurs des câbles à changer lors d'un futur remplacement. (EDMS No ) RELEASED Installation de nouveaux chemins de câbles en LSS1+ et LSS2- du SPS. (EDMS No v.2) IN WORK TPSN replacement in LSS2 of SPS. (EDMS No v.1) UNDER APPROVAL

19 Access rights and safety equipment
Everyone who enters the SPS MUST have the following: Completed the correct courses. Safety course levels 1,2 and 3. Habilitation Electrique. Radiological Protection. Self Rescue Mask Training. The correct access authorisation (via EDH). A valid approved IMPACT. (with their name included on this IMPACT) They must know the number of this IMPACT. Safety shoes. Biocell. Electronic dosimetre. Passive dosimetre. Helmet plus Light. Please note that the medical service no longer contact you when your medical is about to expire!!!!!!! You will receive an 2 months before informing you that some of your access rights will soon expire. It is up to you to contact the medical service to arrange a medical.

20 Important information
Useful Contact Numbers SPS Coordination Team David MCFARLANE Franck BAIS – James RIDEWOOD Control Room SPS CCC – 77500 TI CCC – 72201 CSA RP Christophe Tromel – Angelito Herve – Nadine Conan (BA1) – Elodie Aubert Kurt Weiss (BA1) Coordination Safety Emmanuel Paulat – John Etheridge

21 A.O.B. The new date for the test AUTOTRANSFERT will probably be on 30th and 31st of August. This work should be transparent and not affect access into the SPS. There are currently 38 IMPACTs that are “LATE”. If you have completed your work then you need to close the related IMPACT. Franck and myself have moved office. We are now in B10 (same phone numbers) I will be absent from CERN from Monday the 5th August until Monday the 26th August. You can contact Franck or James on matter concerning the SPS. TT10 – James Ridewood LSS1+ - Sven De Man/ Nadine Conan TSO/IMPACT related - Franck Baïs

22 SPS coordination website
\\cernhomeJ.cern.ch\J\jocksoft\Public\LS1 Next meeting 28/08/2013 892-1-D20


Download ppt "SPS LS1 Planning meeting (N018) 31/07/13"

Similar presentations


Ads by Google